OpenStudy (anonymous):

f(x)=2x+e^(2x). find derivative f-1=(4+e^4). note that f(2)=4+e^4

OpenStudy (turingtest):

this question is poorly typed, I do not understand f-1=(4+e^4) please edit it so it is understandable

OpenStudy (anonymous):

inverse function

OpenStudy (turingtest):

you're saying the inverse function is a constant f(2) ?

OpenStudy (anonymous):

\[f ^{-1} (4+e ^{4})\]

OpenStudy (anonymous):

help!!! please math!!!!! due in 5 minutes

OpenStudy (turingtest):

...cuz that's watchya wrote

OpenStudy (anonymous):

that's the question and i don't understand it myself :/

OpenStudy (turingtest):

wait, is it\[f^{-1}(4+e^4)\]or\[f^{-1}=(4+e^4)\]

OpenStudy (anonymous):

Let f(x)= \[2x+e ^{2x}\]. Find d/dx (\[(f ^{-1}(4+e ^{4}))\]

OpenStudy (anonymous):

Note that f(2)=\[4+e ^{4}\]

OpenStudy (turingtest):

ok, hm.... sorta odd Q gotta think about it

OpenStudy (turingtest):

oh this is easy, we just need this one formula....

OpenStudy (anonymous):

i spent 10 mins trying to get started and couldn't even solve it. You're a super genius :)

OpenStudy (anonymous):

and a life saver. i'm studying for my upcoming tests.

OpenStudy (turingtest):

\[[f^{-1}](a)={1\over f'(f^{-1}(a))}\]and thanks, but I'm no genius, just someone who has practiced :)

OpenStudy (turingtest):

whoops, typo\[[f^{-1}]'(a)={1\over f'(f^{-1}(a))}\]

OpenStudy (anonymous):

i practice and still can't get things right :(

OpenStudy (turingtest):

give it time ;) anyway do you see any ways to use this formula yet?

OpenStudy (anonymous):

yes. i can find the derivative of the f(x) function

OpenStudy (turingtest):

well what is \(a\) in our case?

OpenStudy (turingtest):

and yes to your above comment :)

OpenStudy (anonymous):

a=f(2)

OpenStudy (anonymous):

which is 4+e^4

OpenStudy (turingtest):

excellent ! keep going

OpenStudy (anonymous):

derivative of f(x)= 2(1+e^2x)

OpenStudy (turingtest):

what is \[f^{-1}(a)\]?

OpenStudy (anonymous):

2

OpenStudy (turingtest):

so now we can just plug into the formula, right?

OpenStudy (turingtest):

\[[f^{-1}]'(a)={1\over f'(f^{-1}(a))}\]

OpenStudy (turingtest):

just for reminders...

OpenStudy (anonymous):

\[1/2+2e ^{4}\]

OpenStudy (turingtest):

that whole part after the 1 is in the denominator, right?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

then good job :) congrads!

OpenStudy (anonymous):

:) thanks! couldn't have done it without your help

OpenStudy (turingtest):

you did most of it yourself ;) I just posted the formula pretty much